C = 450 + 0.4Y
I = 350

G = 150
X = 70
Z = 35 + 0.1Y
T = 0.15Y
Yf = 1550



Q.2.4 Calculate the tax revenue to the government of this country when the economy
remains in equilibrium.
(2)
1
Expert's answer
2020-12-03T07:05:14-0500

Y = C + I + G + X - IM

Y = 450 + 0.4Y + 350 + 150 + 70 + 35 + 0.1Y + 0.15Y + 1550

Y - 0.4Y - 0.1Y = 2605

0.5Y/0.5 = 2605/0.5

Y = 5210

Therefore,

Tax revenue will be = 5210 - 1550

= 3660